# SLiM: One-shot Quantization and Sparsity with Low-rank Approximation for LLM Weight Compression

### v2

@@ 5 Conclusion
6 Acknowledgments
@@ Appendix A Notations
SLiM-Naive A one-shot low-rank adapter that minimizes the norm of the difference between the original and the compressed weights.
@@
Appendix B Additional sparse-only results
@@
Appendix C Additional quantization-only results
@@
SLiM-Quant - 31.98 36.46 36.19 40.08 45.61 38.27 31.11 30.51
SLiM-Quant SLiM-Naive 35.29 36.02 42.48 45.01 47.75 48.38 55.96 60.85
SLiM-Quant SLiM-LoRA 35.69 36.42 42.59 45.26 48.18 48.52 56.26 60.59
SLiM-Quant SLiM-LoRA + FT 35.91 36.61 43.29 45.58 48.29 49.04 56.51 60.65
Appendix D Additional fine-tuning results
@@
Naive-LoRA SLiM-Quant 34.28 33.38 38.36 41.21 44.91 45.25 48.45 51.94
Naive-LoRA + FT SLiM-Quant 34.41 34.70 39.72 42.88 46.16 46.76 50.89 55.70
SLiM-LoRA SLiM-Quant 34.62 34.36 40.61 42.73 45.99 46.09 51.15 54.94
SLiM-LoRA + FT SLiM-Quant 35.03 34.58 41.11 43.35 46.71 47.25 52.12 56.60
SLiM-LoRAQ SLiM-Quant 34.43 34.30 40.11 42.37 46.33 46.24 51.02 53.55
SLiM-LoRAQ + FT SLiM-Quant 34.92 34.85 41.84 43.87 46.31 46.91 48.31 56.50
@@
Naive-LoRA SLiM-Quant 34.77 34.23 40.40 43.37 46.64 47.30 51.52 55.33
Naive-LoRA + FT SLiM-Quant 35.70 35.47 41.89 44.16 47.08 47.78 52.90 57.08
SLiM-LoRA SLiM-Quant 35.20 35.32 41.85 43.48 47.08 47.96 54.26 57.85
SLiM-LoRA + FT SLiM-Quant 35.59 35.71 42.37 44.58 47.69 48.26 54.69 57.96
SLiM-LoRAQ SLiM-Quant 35.35 35.13 41.74 43.63 47.16 47.86 54.18 57.33
SLiM-LoRAQ + FT SLiM-Quant 35.65 35.67 42.74 44.54 47.48 48.40 53.57 57.78
Appendix E Language modeling experiments
We evaluate all benchmarks from Section 4 and Appendix B and C on the WikiText2 language modeling task. Tables 8 and 9 show perplexity results for 4-bit quantized models with 2:4 and unstructured sparsity, respectively. To examine sparsity and quantization independently, Tables 10 and 11 report results for pruning-only and quantization-only models. Consistent with Section 4, SLiM achieves superior performance across all settings.
@@ Appendix E Language modeling experiments
SparseGPT Group OPTQ 78.18 59.86 27.36 18.62 15.31 13.25 15.01 8.97
Wanda Group AbsMax 1.8E2 1.3E2 32.76 24.48 17.29 16.86 13.46 8.70
@@
Naive-LoRA SLiM-Quant 83.08 58.69 27.06 20.92 14.29 13.20 8.19 7.09
Naive-LoRA + FT SLiM-Quant 51.82 38.84 20.59 16.19 13.13 12.55 6.96 6.01
SLiM-LoRA SLiM-Quant 57.91 50.09 19.64 15.65 12.71 12.13 7.56 6.50
SLiM-LoRA + FT SLiM-Quant 44.03 37.32 18.25 14.89 12.68 12.06 6.70 6.60
SLiM-LoRAQ SLiM-Quant 53.09 46.96 19.62 16.01 12.48 12.15 7.75 6.96
SLiM-LoRAQ + FT SLiM-Quant 42.80 37.39 18.38 15.40 12.65 12.35 7.08 6.36
@@
Naive-LoRA SLiM-Quant 46.66 33.90 19.46 15.36 12.16 11.41 6.56 5.58
Naive-LoRA + FT SLiM-Quant 38.05 29.27 17.52 14.39 12.28 11.84 6.10 5.28
SLiM-LoRA SLiM-Quant 39.62 31.51 16.52 13.65 11.42 10.82 6.16 5.36
SLiM-LoRA + FT SLiM-Quant 34.92 28.67 16.16 13.66 11.83 11.47 5.36 5.19
SLiM-LoRAQ SLiM-Quant 38.79 30.16 16.64 13.82 11.43 10.80 6.26 5.58
SLiM-LoRAQ + FT SLiM-Quant 35.17 28.31 16.46 13.96 11.42 10.80 5.94 5.46
@@
SLiM-Quant - 1.4E3 26.0 1.7E3 33.1 31.0 6.7E2 1.3E5 7.8E4
SLiM-Quant Naive-LoRA 32.1 24.1 15.6 13.4 11.2 10.5 5.4 4.8
SLiM-Quant SLiM-LoRA 30.8 23.1 15.2 12.9 11.1 10.3 5.4 4.8
SLiM-Quant SLiM-LoRA + FT 30.7 23.5 15.3 13.3 11.6 10.0 5.3 4.7
Appendix F Additional Sparse and Quantized Results
@@
Appendix G Sparsity vs. quantization
A natural question that arises compressing models is whether it is more efficient to reduce the model size through pruning or quantization. To answer this question, we conduct a set of experiments, which evaluate the perplexity of different models under three different conditions, all with around 8×8$\times$8 × model size reduction factor: (1) 2-bit weight quantization with no sparsity, (2) 4-bit weight quantization with 50% unstructured sparsity, and (3) 4-bit weight quantization with 50% 2:4 sparsity. We have used SLiM-LoRA with SLiM-Quant in all the experiments. The accuracy and perplexity results of these experiments are summarized in Tables 13 and 14, showing that combining sparsity and quantization yields better results in comparison to quantization-only settings with lower bitwidth.
@@
Appendix H Additional speedup results
Section 4 presents the speedup of SLiM on consumer-grade GPUs, while this section provides results on NVIDIA A100-40GB GPUs. Figure 3 summarizes the speedup for the LLaMA-2 and LLaMA-3.1 model families, including LLaMA-3.1-405B, highlighting SLiM ’ scalability to large models. As with consumer-grade devices, larger models achieve higher speedups. However, for smaller models like LLaMA-2-7B, Sparse Marlin’s sparse quantized matrix multiplication kernels lead to a slowdown on A100 GPUs, which does not occur on consumer-grade GPUs and is not specific to SLiM .
Figure 3: SLiM speedup for LLaMA-2 family of models on NVIDIA A100-40GB GPUs.
Appendix I Fine-tuning costs
@@
SLiM-Naive SLiM-Quant 1.5h 3h 6h 8h 16h 18h 14h 14h
SLiM-LoRA SLiM-Quant
Appendix J Memory reduction analysis
SLiM prunes and quantizes the models and adds additional low-rank adapters to them. Additionally, it supports quantization methods for the low-rank adapters to reduce their overheads. In the following, we propose an analysis on the reduced memory when using SLiM and other pruning and quantization methods.
@@
Appendix K Computation reduction analysis
@@
Appendix L Compression costs
@@
Appendix M Rank analysis
@@
Appendix N Effects of calibration sample count
@@
Appendix O Sensitivity to calibration dataset
@@
Appendix P Sparsity analysis
@@
Appendix Q Related work
@@
Q.2 Quantization
Quantizing all elements in a matrix is challenging due to the significant impact of outliers on the model (Dettmers et al., 2022). Group quantization (Alistarh et al., 2017; Gunho et al., 2022) addresses this by quantizing small groups of a weight matrix with a shared quantization parameter, but it introduces challenges discussed in Appendix S.
@@
Advanced methods like JSQ (Guo et al., 2024) jointly prune and quantize weights to 8 bits but struggle to recover accuracy in low bit-width quantization, limiting their utility.
Q.3 Low-rank Adapters
Low-rank adapters were first introduced to LLMs to reduce the overhead of fine-tuning (Hu et al., 2021; Mozaffari et al., 2024). Q-LoRA (Dettmers et al., 2023) extended this approach by quantizing weights before fine-tuning, allowing the process to recover accuracy lost during quantization. LQ-LoRA (Guo et al., 2023) further improved Q-LoRA by initializing the adapters using the SVD of the quantization error. LoSparse (Li et al., 2023) has a similar approach as LQ-LoRA, but for sparsity, initializing the low-rank adapters to the norm of the pruning error. RoSA (Nikdan et al., 2024) expands the learning capability of the model by adding both low-rank and sparse adapters to the model. This approach adds and extra sparse matrix multiplication to the inference, increasing the adapter overhead even further. However, all these methods require hundreds of millions of tokens for fine-tuning, making them costly and not comparable to one-shot pruning and quantization methods, or methods that use much shorter fine-tuning phases.
L2QER (Zhang et al., 2024) avoids fine-tuning by using one-shot low-rank adapters to mitigate quantization error. However, it performs poorly when combined with sparsity, resulting in a significant accuracy gap between the compressed and dense models.
Appendix R Settings and hyperparameters
@@
Appendix S Group quantization challenges
Group quantization allows sharing the same quantization parameters for a small group of the elements in the quantized matrix, leading to smaller error. But, using group quantization adds additional challenges to the training and inference of the model, e.g. more complicated implementation and additional memory and compute overheads.
The state-of-the-art group quantization GPU kernel, dense and sparse Marlin (Frantar et al., 2024), consists of thousands of lines of CUDA code optimized for only a limited number of GPU architectures, showcasing the amount of effort needed to implement a version of group quantization. Furthermore, other libraries frameworks, such as Triton (Tillet et al., 2019) and CUTLASS (NVIDIA Corporation, 2025) do not provide support for 4-bit group quantization, limiting its flexibility and possibility of modification.

### v3

@@ 5 Conclusion
Impact Statement
The SLiM framework advances model compression by enabling efficient, one-shot quantization and sparsity for large language models (LLMs) while maintaining accuracy through low-rank approximation. This has the potential to make LLMs more accessible and sustainable by reducing their computational and energy requirements, thereby enabling deployment on a wider range of devices, such as smartphones and edge computing platforms, and contributing to environmental sustainability. However, the increased accessibility of compressed models raises important considerations regarding potential accuracy trade-offs in critical applications, such as healthcare or legal systems, and the ethical implications of broader AI deployment, including risks of bias propagation and misuse. To address these challenges, it is crucial to ensure that efficiency gains do not compromise model reliability and that appropriate safeguards, such as transparency and rigorous evaluation, are in place for responsible AI development.
Acknowledgments
@@ Appendix A Notations
Naive-LoRA A one-shot low-rank adapter that minimizes the norm of the difference between the original and the compressed weights.
@@
Appendix B Input Quantization
We evaluate SLiM with 8-bit input quantization to assess its impact on accuracy. We use AbsMax uniform quantization with a single parameter per input tensor and apply FP8 format (Micikevicius et al., 2022) for weight quantization. The choice between E4M3 and E5M2 depends on the tensor’s maximum value; if it exceeds E4M3’s range, we switch to E5M2 for greater expressivity. Next, we examine how input quantization affects model accuracy.
Table 5 presents accuracy results for different SLiM variants with input quantization. A comparison with Table 9, which reports accuracy without input quantization, reveals minimal accuracy loss, demonstrating SLiM’s robustness. For further validation, we extend these experiments to language modeling tasks (Appendix G).
Table 5: Average zero-shot accuracy of LLaMA-2 and OPT models with 4-bit weight quantization and 8-bit input quantization with 50% weight sparsity. $\uparrow$ indicates better performance.
Pruning/LoRA Weight OPT LLaMA-2
Method Quantization 125M 350M 1.3B 2.7B 6.7B 13B 7B 13B
Dense - 35.9 37.1 43.4 45.5 48.3 48.7 56.6 60.8
50% 2:4
SLiM-LoRA SLiM-QuantW 34.85 34.27 40.29 42.58 45.78 46.21 50.99 54.66
SLiM-LoRA + FT SLiM-QuantW 35.28 34.33 41.14 43.29 46.44 47.33 51.77 56.28
SLiM-LoRAQ SLiM-QuantW 34.30 33.85 39.92 41.99 46.08 45.94 50.70 53.56
SLiM-LoRAQ + FT SLiM-QuantW 34.92 34.80 41.66 43.69 46.03 46.87 50.26 56.28
50% Unstructured
SLiM-LoRA SLiM-QuantW 35.12 34.86 41.94 43.53 47.27 47.70 54.28 57.82
SLiM-LoRA + FT SLiM-QuantW 35.18 35.30 42.37 44.02 47.01 48.52 54.43 57.70
SLiM-LoRAQ SLiM-QuantW 35.26 34.67 41.48 43.46 47.25 47.76 53.91 57.16
SLiM-LoRAQ + FT SLiM-QuantW 35.52 35.31 42.66 44.50 47.08 48.53 53.23 57.55
Appendix C SLiM-QuantW vs. SLiM-QuantO
Table 6 compares the average accuracy of different models when using SLiM-Quant with weight error minimization (SLiM-QuantW) and activation-aware output error minimization (SLiM-QuantO). SLiM-QuantO outperforms SLiM-QuantW by a small gap, while adding computational overhead with irregular memory access patterns at inference time.
Table 6: Average zero-shot accuracy of LLaMA-2 and OPT models with 50% sparsity and 4-bit weight quantization for SLiM-QuantW and SLiM-QuantO.
Pruning/LoRA Weight OPT LLaMA-2
Method Quantization 125M 350M 1.3B 2.7B 6.7B 13B 7B 13B
Dense - 35.9 37.1 43.4 45.5 48.3 48.7 56.6 60.8
2:4 Sparsity
SLiM-LoRA SLiM-QuantW 34.62 34.36 40.61 42.73 45.99 46.09 51.15 54.94
SLiM-LoRA SLiM-QuantO 34.63 34.36 40.29 42.45 45.71 46.24 51.22 55.05
50% Unstructured
SLiM-LoRA SLiM-QuantW 35.20 35.32 41.85 43.48 47.08 47.96 54.26 57.85
SLiM-LoRA SLiM-QuantO 35.20 34.78 41.29 43.31 47.09 47.86 54.46 57.97
Appendix D Additional Sparse-only Results
@@
Appendix E Additional Quantization-only Results
@@
SLiM-QuantW - 31.98 36.46 36.19 40.08 45.61 38.27 31.11 30.51
SLiM-QuantW SLiM-Naive 35.29 36.02 42.48 45.01 47.75 48.38 55.96 60.85
SLiM-QuantW SLiM-LoRA 35.69 36.42 42.59 45.26 48.18 48.52 56.26 60.59
SLiM-QuantW SLiM-LoRA + FT 35.91 36.61 43.29 45.58 48.29 49.04 56.51 60.65
Appendix F Additional Fine-tuning Results
@@
Naive-LoRA SLiM-QuantW 34.28 33.38 38.36 41.21 44.91 45.25 48.45 51.94
Naive-LoRA + FT SLiM-QuantW 34.41 34.70 39.72 42.88 46.16 46.76 50.89 55.70
SLiM-LoRA SLiM-QuantW 34.62 34.36 40.61 42.73 45.99 46.09 51.15 54.94
SLiM-LoRA + FT SLiM-QuantW 35.03 34.58 41.11 43.35 46.71 47.25 52.12 56.60
SLiM-LoRAQ SLiM-QuantW 34.43 34.30 40.11 42.37 46.33 46.24 51.02 53.55
SLiM-LoRAQ + FT SLiM-QuantW 34.92 34.85 41.84 43.87 46.31 46.91 48.31 56.50
@@
Naive-LoRA SLiM-QuantW 34.77 34.23 40.40 43.37 46.64 47.30 51.52 55.33
Naive-LoRA + FT SLiM-QuantW 35.70 35.47 41.89 44.16 47.08 47.78 52.90 57.08
SLiM-LoRA SLiM-QuantW 35.20 35.32 41.85 43.48 47.08 47.96 54.26 57.85
SLiM-LoRA + FT SLiM-QuantW 35.59 35.71 42.37 44.58 47.69 48.26 54.69 57.96
SLiM-LoRAQ SLiM-QuantW 35.35 35.13 41.74 43.63 47.16 47.86 54.18 57.33
SLiM-LoRAQ + FT SLiM-QuantW 35.65 35.67 42.74 44.54 47.48 48.40 53.57 57.78
Appendix G Language Modeling Experiments
We evaluate all benchmarks from Section 4 and Appendix B ,D, and E on the WikiText2 language modeling task. Tables 10 and 11 show perplexity results for 4-bit quantized models with 2:4 and unstructured sparsity, respectively. Table 12 summarizes the results for 8-bit input quantization. To examine sparsity and quantization independently, Tables 13 and 14 report results for pruning-only and quantization-only models. Consistent with Section 4, SLiM achieves superior performance across all settings.
@@ Appendix G Language Modeling Experiments
SparseGPT Group OPTQ 78.18 59.86 27.36 18.62 15.31 13.25 12.07 9.46
Wanda Group AbsMax 1.8E2 1.3E2 32.76 24.48 17.29 16.86 14.36 9.38
@@
Naive-LoRA SLiM-QuantW 83.08 58.69 27.06 20.92 14.29 13.20 8.19 7.09
Naive-LoRA + FT SLiM-QuantW 51.82 38.84 20.59 16.19 13.13 12.55 6.96 6.01
SLiM-LoRA SLiM-QuantW 57.91 50.09 19.64 15.65 12.71 12.13 7.77 6.80
SLiM-LoRA + FT SLiM-QuantW 44.03 37.32 18.25 14.89 12.68 12.06 6.70 6.60
SLiM-LoRAQ SLiM-QuantW 53.09 46.96 19.62 16.01 12.48 12.15 7.75 6.96
SLiM-LoRAQ + FT SLiM-QuantW 42.80 37.39 18.38 15.40 12.65 12.35 7.08 6.36
@@
Naive-LoRA SLiM-QuantW 46.66 33.90 19.46 15.36 12.16 11.41 6.56 5.58
Naive-LoRA + FT SLiM-QuantW 38.05 29.27 17.52 14.39 12.28 11.84 6.10 5.28
SLiM-LoRA SLiM-QuantW 39.62 31.51 16.52 13.65 11.42 10.82 6.16 5.36
SLiM-LoRA + FT SLiM-QuantW 34.92 28.67 16.16 13.66 11.83 11.47 5.36 5.19
SLiM-LoRAQ SLiM-QuantW 38.79 30.16 16.64 13.82 11.43 10.80 6.26 5.58
SLiM-LoRAQ + FT SLiM-QuantW 35.17 28.31 16.46 13.96 11.42 10.80 5.94 5.46
Table 12: Perplexity of LLaMA-2 and OPT models with 4-bit weight quantization and 8-bit input quantization. $\downarrow$ indicates better performance.
Pruning/LoRA Weight OPT LLaMA-2
Method Quantization 125M 350M 1.3B 2.7B 6.7B 13B 7B 13B
Dense - 27.66 22.00 14.62 12.47 10.86 10.13 5.47 4.89
50% 2:4
SLiM-LoRA SLiM-QuantW 48.4 49.6 16.6 16.2 12.9 12.3 7.2 6.5
SLiM-LoRA + FT SLiM-QuantW 39.8 37.5 18.3 15.5 12.8 12.1 6.6 5.8
SLiM-LoRAQ SLiM-QuantW 54.2 50.8 20.8 16.8 13.0 12.4 7.8 7.0
SLiM-LoRAQ + FT SLiM-QuantW 43.4 39.1 19.3 16.0 13.1 12.6 7.1 5.8
50% Unstructured
SLiM-LoRA SLiM-QuantW 36.8 31.1 16.8 14.0 11.7 10.9 6.1 5.4
SLiM-LoRA + FT SLiM-QuantW 33.8 28.6 16.5 14.0 12.0 11.5 5.9 5.2
SLiM-LoRAQ SLiM-QuantW 39.5 31.3 17.3 14.2 11.8 10.9 6.3 5.6
SLiM-LoRAQ + FT SLiM-QuantW 35.6 29.1 17.0 14.3 12.2 11.7 6.2 5.5
@@
SLiM-QuantW - 1.4E3 26.0 1.7E3 33.1 31.0 6.7E2 1.3E5 7.8E4
SLiM-QuantW Naive-LoRA 32.1 24.1 15.6 13.4 11.2 10.5 5.4 4.8
SLiM-QuantW SLiM-LoRA 30.8 23.1 15.2 12.9 11.1 10.3 5.4 4.8
SLiM-QuantW SLiM-LoRA + FT 30.7 23.5 15.3 13.3 11.6 10.0 5.3 4.7
Appendix H Additional Sparse and Quantized Results
@@
Appendix I Sparsity vs. Quantization
A natural question that arises compressing models is whether it is more efficient to reduce the model size through pruning or quantization. To answer this question, we conduct a set of experiments, which evaluate the perplexity of different models under three different conditions, all with around 8×8$\times$8 × model size reduction factor: (1) 2-bit weight quantization with no sparsity, (2) 4-bit weight quantization with 50% unstructured sparsity, and (3) 4-bit weight quantization with 50% 2:4 sparsity. We have used SLiM-LoRA with SLiM-Quant in all the experiments. The accuracy and perplexity results of these experiments are summarized in Tables 16 and 17, showing that combining sparsity and quantization yields better results in comparison to quantization-only settings with lower bitwidth.
@@
Appendix J Additional Speedup Results
Section 4 presents the speedup of SLiM on consumer-grade GPUs, while this section provides results on NVIDIA A100-40GB GPUs. Figure 4 summarizes the speedup for the LLaMA-2 and LLaMA-3.1 model families, including LLaMA-3.1-405B, highlighting SLiM ’ scalability to large models. As with consumer-grade devices, larger models achieve higher speedups.
Additionally, the breakdown of the speedup, showing the contribution of quantization and sparsity, is demonstrated using brighter and darker colors, respectively.
Figure 4: SLiM speedup for LLaMA-2 family of models on NVIDIA A100-40GB GPUs. The brighter color shows the contribution of quantization to the total speedup.
Appendix K Fine-tuning Costs
@@
SLiM-Naive SLiM-QuantW 1.5h 3h 6h 8h 16h 18h 14h 14h
SLiM-LoRA SLiM-QuantW
Appendix L Memory Reduction Analysis
SLiM prunes and quantizes the models and adds additional low-rank adapters to them. Additionally, it supports quantization methods for the low-rank adapters to reduce their overhead. In the following, we propose an analysis of the reduced memory when using SLiM and other pruning and quantization methods.
@@
Appendix M Computation Reduction Analysis
@@
Appendix N Compression Costs
@@
Appendix O Rank Analysis
@@
Appendix P Effects of Calibration Sample Count
@@
Appendix Q Sensitivity to Calibration Dataset
@@
Appendix R Sparsity Analysis
@@
Appendix S Related Work
@@
In addition to post-training sparsity, a recent line of work targets sparsity during training (Lu et al., 2023; Mozaffari et al., 2024; Bambhaniya et al., 2024); however, their applicability is limited because of the expensive costs of training.
S.2 Quantization
Quantizing all elements in a matrix is challenging due to the significant impact of outliers on the model (Dettmers et al., 2022). Group quantization (Alistarh et al., 2017; Gunho et al., 2022) addresses this by quantizing small groups of a weight matrix with a shared quantization parameter, but it introduces challenges discussed in Appendix U.
@@
Advanced methods like JSQ (Guo et al., 2024) jointly prune and quantize weights to 8 bits but struggle to recover accuracy in low bit-width quantization, limiting their utility. An analysis of the interplay between sparsity and quantization can be found in (Harma et al., 2024).
S.3 Low-rank Adapters
Low-rank adapters were first introduced to LLMs to reduce the overhead of fine-tuning (Hu et al., 2021; Mozaffari et al., 2024). Q-LoRA (Dettmers et al., 2023) extended this approach by quantizing weights before fine-tuning, allowing the process to recover accuracy lost during quantization. LQ-LoRA (Guo et al., 2023) further improved Q-LoRA by initializing the adapters using the SVD of the quantization error. LoSparse (Li et al., 2023) has a similar approach as LQ-LoRA, but for sparsity, initializing the low-rank adapters to the norm of the pruning error. RoSA (Nikdan et al., 2024) expands the learning capability of the model by adding both low-rank and sparse adapters to the model. This approach adds an extra sparse matrix multiplication to the inference, increasing the adapter overhead even further. However, all these methods require hundreds of millions of tokens for fine-tuning, making them costly and not comparable to one-shot pruning and quantization methods, or methods that use much shorter fine-tuning phases.
L2QER (Zhang et al., 2024a) avoids fine-tuning by using one-shot low-rank adapters to mitigate quantization error. However, it performs poorly when combined with sparsity, resulting in a significant accuracy gap between the compressed and dense models. More recent methods, QERA (Zhang et al., 2024b) and CALDERA (Saha et al., 2024) find closed-form solutions to the problem discussed in L2QER, but they still do not support sparsity.
Appendix T Settings and Hyperparameters
@@
Appendix U Group Quantization Challenges
Group quantization allows sharing the same quantization parameters for a small group of the elements in the quantized matrix, leading to smaller errors. But, using group quantization adds additional challenges to the training and inference of the model, e.g. more complicated implementation and additional memory and compute overheads.
The state-of-the-art group quantization GPU kernel, dense and sparse Marlin (Frantar et al., 2024), consists of thousands of lines of CUDA code optimized for only a limited number of GPU architectures, showcasing the amount of effort needed to implement a version of group quantization. Furthermore, other libraries and frameworks, such as Triton (Tillet et al., 2019) and CUTLASS (NVIDIA Corporation, 2025) do not provide support for 4-bit group quantization, limiting its flexibility and possibility of modification.